OpenStudy (anonymous):

Which of the following statements have the same result? Explain each step in solving each one. f(1) when f(x) = 5x + 1 f−1(3) when f(x) = 2x+3 3y − 7 = y + 5

OpenStudy (anonymous):

@Kim21 @jim_thompson5910

OpenStudy (anonymous):

so you have 1/ f(2) = 8 let it there now, 2/ f−1(3) when f(x)=2x−73 you have to find f−1(x) which is 3x+72 and then f−1(3)=8 So, 1/ = 2 2/ = 8 now, 3/ solve for y : 2y+14 = 4y -2 16 = 2y y = 8 Summary : 1/ = 2 2/ = 8 3/ =8 conclusion: 2/ and 3/ have the same solution. Read it carefully, hopefully you can follow and know how to work on this kind of problem from now on. Good luck

OpenStudy (anonymous):

oh, mistake, f(2 ) =8 , they are all equal. I am sorry.

OpenStudy (anonymous):

Lol how did you do that so fast?

OpenStudy (anonymous):

i answerd the same question a few days ago :D copy paste lol

OpenStudy (anonymous):

so f(x)5x+1=?

OpenStudy (anonymous):

yayysss<3 Thanks so much!

OpenStudy (anonymous):

welcome :)